A solution has a [OH–] of 2.5 x 10–7M, what colour would it be if it was mixed with the indicator phenol red?
1
Expert's answer
2020-05-19T08:43:28-0400

Solution:



[OH] = 2.5×10–7 M

To calculate the pOH of a solution, use the equation:

pOH = −log[OH]

pOH = −log[2.5×10–7] = 6.60

pOH = 6.60


pH + pOH = 14

pH = 14 − pOH = (14 − 6.60) = 7.4

pH = 7.4